Projectile Motion and Impact: Just for fun, a golf
Answer #1

consider maximum height at a certain bounce number be h.
at maximum height vertical speed =0

time taken to reach ground=sqrt(2*h/g)

velocity in this time=initial velocity+acceleration*time

=0+g*sqrt(2*h/g)=sqrt(2*g*h)

as coefficient of restitution is 0.92,

then the velocity with which the ball will bounce back=0.92*sqrt(2*g*h)

then maximum height reached in next stage=initial velocity^2/(2*g)

=0.92^2*2*g*h/(2*g)=0.92^2*h

using the same analysis,

for nth bounce, maximum height=0.92^(2*n)*h0

part a:

using n=3, h3=0.92^(6)*1.5=0.90953 m

part b:

maximum vertical height after nth bounce=e^(2*n)*h0

part c:

let at nth bounce, bounce height=0.5*42.7 mm=21.35 mm=0.02135 m

then 0.92^(2*n)*1.5=0.02135

==>0.92^(2*n)=0.014233

taking logarith,

2*n=50.996

==>n=25.498

hence after 26th bounce, the ball will start rolling.